# Differentiable Physics-based System Identification for Robotic Manipulation of Elastoplastic Materials

### v1

@@ 1 Introduction
The intrinsic complexity of these dynamics inhibits the direct application of conventional robotic motion planning methods, which typically require explicit physics models for all concerned objects Latombe (2012). Learning approaches, such as reinforcement learning (RL), often involve training an agent to learn to interpret its perception and take actions through inefficient trial & error in a realistic physics-based simulation Collins et al. (2021); Kroemer et al. (2021), which is both challenging and largely unavailable when it comes to deformable materials.
@@
In contrast to the well-studied rigid body dynamics in robotics, where motions can be predicted and controlled using well-defined equations of motion and deterministic models Featherstone (2014), deformable materials do not follow such straightforward patterns. Directly applying these methods is problematic because it is extremely difficult and often infeasible to accurately model and perceive real-world elastoplastic materials and measure the underlying physics parameters that govern their motions and deformations  Arriola-Rios et al. (2020); Yin et al. (2021). As a result, achieving high-precision manipulation for such materials with motion planning or data-driven techniques is challenging due to the high computational cost and the lack of techniques to capture the dynamics accurately.
@@
To close this gap, this research proposes a Differentiable Physics-based System Identification (DPSI) framework for the robotic manipulation of volumetric elastoplastic materials. Our framework can efficiently estimate key physics parameters governing material deformation dynamics using minimal and simple manipulation motions. The estimated physics parameters enable accurate material simulation for long-horizon predictions of real-world elastoplastic deformation behaviours.
@@
The workflow of the proposed DPSI framework can be summarised as follows. As shown in Figure 1, a robot equipped with an in-hand 3D camera (Zivid) and three end-effectors is deployed to manipulate the elastoplastic object (e.g., play dough). Before manipulation, the robot takes multi-view point clouds of the object to minimise occlusions. These point clouds are used to create the initial particle system for the material point methods (MPM)-based simulation. The robot then performs a manipulation motion on the object and captures point clouds of the deformed state. Physics simulations of the same manipulation are run with the same initial state and motion, whose resultant particle states are compared to the real-world deformed state through variants of the Chamfer distance (CD) and the earth mover’s distance (EMD) loss. The parameters of the simulated physics models are updated after every simulation to minimise the loss. To facilitate fast optimisation, we use a differentiable simulator built on the TaiChi auto-differentiation mechanism, which allows automatic gradient computation from the losses and gradient-based optimisation for the physics parameters Hu et al. (2019, 2020).
@@
High-fidelity physics: Unlike previous methods that either employ non-physics-based models (e.g., neural networks) or highly simplified material geometry representations (e.g., sparse keypoints), we use high-fidelity physics-based simulation powered by the MPM Jiang et al. (2016), which simulates materials as Lagrangian particles and keeps track of their positions and velocities. It achieves faster simulation by computing the motions, deformation gradients, and frictional contacts on a background Eulerian grid Stomakhin et al. (2013); Jiang et al. (2016); Gao et al. (2017); Hu et al. (2018). MPM-based simulations provide highly efficient and realistic simulation with high physical plausibility by closely following real-world physics laws such as Newton’s laws and elastic and plastic energy conservation models.
@@
Incomplete & noisy observations: Unlike existing works that rely on synthetic videos with complete sequences of perfect observations Murthy et al. (2020); Li et al. (2023); Kaneko (2024), our framework uses 3D point clouds to observe real-world object geometries. Capturing the full depth of an object during manipulation is impractical due to occlusions caused by the end-effector or environment. This means that only the point clouds before and after a manipulation motion are practical to obtain and informative enough to observe the full geometry of the deformed object in real-world experiments. In addition, real-world point clouds tend to suffer from inaccurately estimated camera matrices and sensory noises.
@@
Small data, short & simple motions: Using extensive and diverse manipulation motions to collect real-world deformation data is time-consuming and costly. Existing studies demand a significant number of real-world interactions or complete sequences of simulation videos to identify object deformations under various motions, yet still resulting in simulations with insufficient accuracy for real-world applications Lin et al. (2022); Shi et al. (2023); Li et al. (2023); Kaneko (2024). Our goal is to recover physics parameters that enable accurate predictions of long-horizon, unseen and complicated elastoplastic material manipulation dynamics, using minimal simple and short real-world interactions.
@@
Joint parameter estimation: We aim to jointly estimate the physics parameters provided by physics models. Besides Newton’s laws, we employ the fixed corotated elastic energy model Stomakhin et al. (2012), the von Mises plasticity model Jones (2009) and the dynamic friction model in our simulation. These lead to six key parameters: Young’s modulus $E$, Poisson’s ratio $\nu$, yield stress $\boldsymbol{\sigma}_{y}$, material density $\rho$, and the friction coefficients of the table $\eta_{t}$ and end-effectors $\eta_{m}$ (assuming the three end-effectors share the same coefficient). The first four parameters primarily govern the deformation responses, while the last two handle frictional contacts. These parameters are heavily intertwined in governing the behaviours of the manipulated object and there is no intuitive solution to identify one of them without estimating the effects of other parameters. Therefore, we seek to identify these parameters simultaneously.
@@
Differentiable physics: Identifying the physics parameters in their discretised spaces via search or evolutionary algorithms is computationally slow due to the exponentially growing number of possible combinations as the discretisation becomes finer. While gradient-based optimisation methods offer faster convergence toward the minimum, it is infeasible with most physics simulations because many computation steps are not differentiable and these simulators do not support derivative computations. In this work, we explore the feasibility of optimising system parameters using gradients computed by differentiating loss functions through a physics simulator written by DiffTaiChi, a programming language tailored for GPU-accelerated parallel computation and automatic differentiation Hu et al. (2019, 2020). DiffTaichi generates derivative functions for simulation steps via source code transformation that retains arithmetic intensity and parallelism. It uses a memory-efficient tape to record the order of computation kernels for forward simulation and traverses their derivative functions in the backward order to generate gradients through the computation graph. We build DPSI upon DiffTaiChi and explore the feasibility of directly optimising several physics parameters jointly with gradients computed by differentiating point-cloud-based loss functions through the high-fidelity physics simulator.
@@
Substantial experiments demonstrate that our main contribution, DPSI, can achieve highly accurate simulation-to-reality alignment for elastoplastic materials manipulated by unseen, long horizon and complex motions using minimal simple and short interactions, and noisy and incomplete observations. Results show that when multiple solutions and parameter uncertainty exist, DPSI can provide physically intuitive parameter interpretations that can guide further system identification, model improvement, and motion adaptation. Statistics on the computation costs of DPSI indicates promising practical deployment of the DPSI framework.
@@ 2.1 Deformable object manipulation
Both model-free and model-based approaches have been taken for manipulating deformable materials. Existing model-free methods often lack manipulation precision due to the absence of physics laws that describe the motion and deformation under complex contacts McConachie et al. (2020); Cherubini et al. (2020); Shi et al. (2023, 2024); Shen et al. (2024). Physics-model-based methods, while more accurate, struggle with aligning the simulated dynamics with the real world and often rely on simplified geometric representations for higher computational efficiency, sacrificing manipulation precision Navarro-Alarcon et al. (2016); Yang et al. (2023); Shetab-Bushehri et al. (2023). With or without a model, many of them use simulations that offer a cost-effective way to test manipulation methods and allow the collection of massive data for learning-based approaches Arriola-Rios et al. (2020); Collins et al. (2021); Yin et al. (2021). However, inaccurate dynamics predictions and control precision are unacceptable for many tasks such as surgery, assembly and disassembly. Therefore, akin to the way humans efficiently learn about object and environment physics properties, this work proposes to actively identify system parameters that enable high-precision simulation of real-world manipulation dynamics for volumetric elastoplastic objects.
@@ 2.2 Deformable object modelling
Simulators are essential for advancing robotic manipulation, providing a fast, low-cost alternative to real-world testing Featherstone (2014); Collins et al. (2021). In recent years, arguably the most efficient and accurate simulation for 3D deformable objects is achieved by the material point method (MPM) Jiang et al. (2016). Like most well-known methods for simulating continuum matter, such as position-based dynamics (PBD) and smoothed particle hydrodynamics (SPH) Yin et al. (2021), MPM represents the object as Lagrangian particles and keeps track of their positions and velocities. Unlike pure Lagrangian methods, MPM achieves faster simulation by computing the motions, deformation gradients, and frictional contacts on a background Eulerian grid (akin to the finite element method), governed by elastic and plastic energy functions and Newton’s laws in the form of partial differential equations Stomakhin et al. (2013); Jiang et al. (2016); Gao et al. (2017); Hu et al. (2018). MPM has been proven superior to other methods in terms of efficiency and visual effects for objects that undergo large deformations, fractures, and self-collision Hu et al. (2018). Despite its advantages, accurately identifying physics parameters for real-world robotic manipulation of elastoplastic objects remains an open challenge.
@@ 2.3 Deformable object system identification
Unlike rigid bodies Jaques et al. (2022); Heiden et al. (2022); Chen et al. (2022), identifying the physics properties of deformable objects is more complex than for rigid bodies, primarily due to the difficulty in measuring key parameters like material properties and friction coefficients Arriola-Rios et al. (2020); Yin et al. (2021). Previous research has mostly focused on linear and planar deformable objects, such as ropes and cloths Sundaresan et al. (2022); Yang et al. (2022); Caporali et al. (2024).
Early works that sought to identify 3D volumetric material properties have much simpler assumptions and scenarios. For instance, one of the earlier works uses gradient-based optimisation to retrieve the stiffness of a spring system representing elastic deformable objects Lloyd et al. (2007), while another work uses an exhaustive search method to find the value of Poisson’s ratio for an elastic form object Güler et al. (2017). They focus on single parameter identification for volumetric elastic deformable objects with reduced Dofs that under-represent the geometries and deformation behaviours of real-life objects.
Recent efforts like GradSim Murthy et al. (2020) have used differentiable physics and rendering to identify five parameters for elastic objects with much higher DoFs from single-view simulation videos, demonstrating the feasibility of differentiable system identification with synthetic videos. Two following works, PAC-NeRF Li et al. (2023) and LPO Kaneko (2024) propose to jointly reconstruct object geometries (shapes, positions and colours) and physical properties by using a voxel neural radiance field Sun et al. (2022) that performs differentiable rendering and allows gradients to be back-propagated from the image space to the Eulerian grid.
Compared to these works that focused on simulations, our study tackles the more challenging task of system identification for real-world objects through robot interactions. Similarly, we also employ the material point method Hu et al. (2018) and DiffTaiChi Hu et al. (2020) for differentiable physics simulation and study the system identification task without known object geometries. However, we aim to align simulations with real-world dynamics with minimal robot interactions, using only incomplete, occluded, and noisy point cloud data.
@@ 3.3 Differentiable dynamics modelling
We modified the standard procedures in one simulation step of the MLS-MPM as shown below to incorporate end-effector control following Xian et al. (2023). The readers are referred to the original paper for more details about MPM Stomakhin et al. (2013) and MLS-MPM Hu et al. (2018).
@@
    Particle to grid. Use the affine particle-in-cell transform Jiang et al. (2015) to transfer the velocities and masses of the Lagrangian particles to the background Eulerian grid nodes. In our implementation, we assume equal volume and mass for the particles.
@@
Elastoplasticity: We assume strain elastoplasticity for the studied objects, meaning that the strain-stress relationship of the material is described using the deformation gradient that can be decomposed into elastic and plastic parts , $\mathbf{F}=\mathbf{F}^{E}\mathbf{F}^{P}$, and the first Piola-Kirchoff stress, $\mathbf{P}$, as $\mathbf{P}=\partial\Psi/\partial\mathbf{F}^{E}$, where $\Psi(\mathbf{F}^{E})$ is the elastic energy density function. In this study, we use the fixed corotated model Stomakhin et al. (2012) as the energy function for its robustness under large deformation, although other constitutive models may be used for different needs. The fixed corotated elastic energy density function is
@@
where $\mathbf{R}^{E}$ is the rotation matrix that comes from the polar decomposition of the deformation gradient, $\mathbf{F}^{E}=\mathbf{R}^{E}\mathbf{S}^{E}$, where $\mathbf{S}^{E}$ is a symmetric matrix. In practice, they are computed from the singular value decomposition of $\mathbf{F}^{E}$ as $\mathbf{F}^{E}=\mathbf{U}^{E}\mathbf{\Sigma}^{E}(\mathbf{V}^{E})^{T}=\mathbf{U}^{E}(\mathbf{V}^{E})^{T}\mathbf{V}^{E}\mathbf{\Sigma}^{E}(\mathbf{V}^{E})^{T}$, and then $\mathbf{R}^{E}=\mathbf{U}^{E}(\mathbf{V}^{E})^{T}$ and $\mathbf{S}^{E}=\mathbf{V}^{E}\mathbf{\Sigma}^{E}(\mathbf{V}^{E})^{T}$ Baker (2005).
@@
In this work, we follow Gao et al. (2017) to use the von Mises model Jones (2009) to compute the return mapping, which takes on the associative plastic flow assumption. The projection process of the trial stress outside of the yield region can be described succinctly as $\tilde{\boldsymbol{\epsilon}}^{E,n}-\boldsymbol{\epsilon}^{E,n}=\delta_{\gamma}\tilde{\boldsymbol{s}}^{E,n}/||\tilde{\boldsymbol{s}}^{E,n}||$, where $\tilde{\boldsymbol{\epsilon}}^{E,n}=\log(\tilde{\mathbf{F}}^{E,n})$ and $\boldsymbol{\epsilon}^{E,n}=\log(\mathbf{F}^{E,n})$ are the trial and modified Hencky strains, $\tilde{\boldsymbol{s}}^{E,n}=\tilde{\boldsymbol{\epsilon}}^{E,n}-\ tr(\tilde{\boldsymbol{\epsilon}}^{E,n})/3\ \boldsymbol{I}$ is the trial deviatoric stress, and $\delta_{\gamma}\tilde{\boldsymbol{s}}^{E,n}/||\tilde{\boldsymbol{s}}^{E,n}||$ is the modifying term based on the solution given in Gao et al. (2017). In particular, based on the von Mises yield criterion, $\sqrt{3J_{2}}-\boldsymbol{\sigma}_{y}\leq 0$, where $\boldsymbol{\sigma}_{y}$ is the yield stress and $J_{2}$ is the second deviatoric stress invariant, it was derived that $\delta_{\gamma}=||\tilde{\boldsymbol{s}}^{E,n}||-\boldsymbol{\sigma}_{y}/2\mu$. With the fact that the singular vectors of the trial elastic deformation gradient do not change, the return mapping operation is done on the eigenvalues of the trial Hencky stress, thus on the eigenvalues of the trial elastic deformation gradient. Then, we can obtain the result of the return mapping by taking the exponential of the modified deviatoric stress: $\Sigma^{E,n}=e^{\boldsymbol{\epsilon}^{E,n}}$ and then the modified deformation gradient: $\mathbf{F}^{E,n}=\tilde{\mathbf{U}}^{E,n}\mathbf{\Sigma}^{E}(\tilde{\mathbf{V}}^{E,n})^{T}$.
@@
Differentiable programming: Several programming tools are available for creating differentiable simulations, such as PyTorch Sundaresan et al. (2022); Arnavaz et al. (2023), DiffTaiChi Hu et al. (2020), Jax Schoenholz and Cubuk (2020) and neural-network-based simulators Heiden et al. (2021). We build our simulator based on DiffTaiChi due to its automated differentiation mechanism, GPU-accelerated parallel computation, fast computation kernel evaluation, intuitive Python APIs, rich community support and various promising applications Huang et al. (2021); Lin et al. (2022); Xian et al. (2023).
In particular, for each numerical computation step, DiffTaiChi flattens its computation branches (e.g., boundary and collisions) and replaces mutable local variables with extra local storage variables, producing straight-line codes without mutable variables. It then uses standard source code transform to generate the derivative function based on the adjoint method. To compute gradients with a loss function, DiffTaiChi records the order of computation kernels and the scalar variables in the forward simulation direction, and then it computes the gradients for the concerned variables by evaluating the derivative functions in the reversed simulation direction Hu et al. (2020). We build a simulator based on this programming language to allow automatic gradient computation for the physics parameters. The codes for the simulator will be open-sourced upon paper acceptance.

### v2

@@ 1 Introduction
The intrinsic complexity of these dynamics inhibits the direct application of conventional robotic motion planning methods, which typically require explicit physics models for all concerned objects (Latombe, 2012). Learning approaches, such as reinforcement learning (RL), often involve training an agent to learn to interpret its perception and take actions through inefficient trial & error in a realistic physics-based simulation (Collins et al., 2021; Kroemer et al., 2021), which is both challenging and largely unavailable when it comes to deformable materials.
@@
In contrast to the well-studied rigid body dynamics in robotics, where motions can be predicted and controlled using well-defined equations of motion and deterministic models (Featherstone, 2014), deformable materials do not follow such straightforward patterns. Directly applying these methods is problematic because it is extremely difficult and often infeasible to accurately model and perceive real-world elastoplastic materials and measure the underlying physics parameters that govern their motions and deformations (Arriola-Rios et al., 2020; Yin et al., 2021). Additional challenges such as time-varying material properties (saturation, dryness, etc.) make the deformation dynamics even more unpredictable. As a result, achieving high-precision manipulation for such materials with motion planning or data-driven techniques is challenging due to the high computational cost and the lack of techniques to capture the dynamics accurately and efficiently.
@@
To close this gap, this research proposes a Differentiable Physics-based System Identification (DPSI) framework for the robotic manipulation of volumetric elastoplastic materials. Our framework can efficiently estimate key physics parameters governing material deformation dynamics using minimal and simple manipulation motions (in less than $5$ minutes with a good initial guess). The estimated physics parameters enable accurate material simulation for long-horizon predictions of real-world elastoplastic deformation behaviours.
@@
The workflow of the proposed DPSI framework can be summarised as follows. As shown in Figure 1, a robot equipped with an in-hand 3D camera (Zivid) and three end-effectors is deployed to manipulate the elastoplastic object (e.g., play dough). Before manipulation, the robot takes multi-view point clouds of the object to minimise occlusions. These point clouds are used to create the initial particle system for the material point methods (MPM)-based simulation. The robot then performs a manipulation motion on the object and captures point clouds of the deformed state. Physics simulations of the same manipulation are run with the same initial state and motion, whose resultant particle states are compared to the real-world deformed state through variants of the Chamfer distance (CD) and the earth mover’s distance (EMD) loss. The parameters of the simulated physics models are updated after every simulation to minimise the loss. To facilitate fast optimisation, we use a differentiable simulator built on the TaiChi auto-differentiation mechanism, which allows automatic gradient computation from the losses and gradient-based optimisation for the physics parameters (Hu et al., 2019, 2020).
@@
High-fidelity physics: Unlike previous methods that either employ non-physics-based models (e.g., neural networks) or highly simplified material geometry representations (e.g., sparse keypoints), we use high-fidelity physics-based simulation powered by the MPM (Jiang et al., 2016), which simulates materials as Lagrangian particles and keeps track of their positions and velocities. It achieves faster simulation by computing the motions, deformation gradients, and frictional contacts on a background Eulerian grid (Stomakhin et al., 2013; Jiang et al., 2016; Gao et al., 2017; Hu et al., 2018). MPM-based simulations provide highly efficient and realistic simulation with high physical plausibility by closely following real-world physics laws such as Newton’s laws and elastic and plastic energy conservation models.
@@
Incomplete & noisy observations: Unlike existing works that rely on synthetic videos with complete sequences of perfect observations (Murthy et al., 2020; Li et al., 2023; Kaneko, 2024), our framework uses 3D point clouds to observe real-world object geometries. Capturing the full depth of an object during manipulation is impractical due to occlusions caused by the end-effector or environment. This means that only the point clouds before and after a manipulation motion are practical to obtain and informative enough to observe the full geometry of the deformed object in real-world experiments. In addition, real-world point clouds tend to suffer from inaccurately estimated camera matrices and sensory noises.
@@
Small data, short & simple motions: Using extensive and diverse manipulation motions to collect real-world deformation data is time-consuming and costly. Existing studies demand a significant number of real-world interactions or complete sequences of simulation videos to identify object deformations under various motions, yet still resulting in simulations with insufficient accuracy for real-world applications (Lin et al., 2022; Shi et al., 2023; Li et al., 2023; Kaneko, 2024). Our goal is to recover physics parameters that enable accurate predictions of long-horizon, unseen and complicated elastoplastic material manipulation dynamics, using minimal simple and short real-world interactions.
@@
Joint parameter estimation: We aim to jointly estimate the physics parameters provided by physics models. Besides Newton’s laws, we employ the fixed corotated elastic energy model (Stomakhin et al., 2012), the von Mises plasticity model (Jones, 2009) and the dynamic friction model in our simulation. These lead to six key parameters: Young’s modulus $E$, Poisson’s ratio $\nu$, yield stress $\boldsymbol{\sigma}_{y}$, material density $\rho$, and the friction coefficients of the table $\eta_{t}$ and end-effectors $\eta_{m}$ (assuming the three end-effectors share the same coefficient). The first four parameters primarily govern the deformation responses, while the last two handle frictional contacts. These parameters are heavily intertwined in governing the behaviours of the manipulated object and there is no intuitive solution to identify one of them without estimating the effects of other parameters. Therefore, we seek to identify these parameters simultaneously.
@@
Differentiable physics: Identifying the physics parameters in their discretised spaces via search or evolutionary algorithms is computationally slow due to the exponentially growing number of possible combinations as the discretisation becomes finer. While gradient-based optimisation methods offer faster convergence toward the minimum, it is infeasible with most physics simulations because many computation steps are not differentiable and these simulators do not support derivative computations. In this work, we explore the feasibility of optimising system parameters using gradients computed by differentiating loss functions through a physics simulator written by DiffTaiChi, a programming language tailored for GPU-accelerated parallel computation and automatic differentiation (Hu et al., 2019, 2020). DiffTaichi generates derivative functions for simulation steps via source code transformation that retains arithmetic intensity and parallelism. It uses a memory-efficient tape to record the order of computation kernels for forward simulation and traverses their derivative functions in the backward order to generate gradients through the computation graph. We build DPSI upon DiffTaiChi and explore the feasibility of directly optimising several physics parameters jointly with gradients computed by differentiating point-cloud-based loss functions through the high-fidelity physics simulator.
@@
Substantial experiments demonstrate that our main contribution, DPSI, can achieve highly accurate simulation-to-reality alignment for elastoplastic materials manipulated by unseen, long horizon and complex motions using minimal simple and short interactions, and noisy and incomplete observations. Results show that when multiple solutions and parameter uncertainty exist, DPSI can provide physically intuitive parameter interpretations that can guide further system identification, model improvement, and motion adaptation. Statistics on the computation costs also indicate promising practical deployment of the DPSI framework.
@@ 2.1 Deformable object manipulation
Both model-free and model-based approaches have been taken for manipulating deformable materials. Existing model-free methods often lack manipulation precision due to the absence of physics laws that describe the motion and deformation under complex contacts (McConachie et al., 2020; Cherubini et al., 2020; Shi et al., 2023, 2024; Shen et al., 2024). Physics-model-based methods, while more accurate, struggle with aligning the simulated dynamics with the real world and often rely on simplified geometric representations for higher computational efficiency, sacrificing manipulation precision (Navarro-Alarcon et al., 2016; Yang et al., 2023; Shetab-Bushehri et al., 2023). With or without a model, many of them use simulations that offer a cost-effective way to test manipulation methods and allow the collection of massive data for learning-based approaches (Arriola-Rios et al., 2020; Collins et al., 2021; Yin et al., 2021). However, inaccurate dynamics predictions and control precision are unacceptable for many tasks such as surgery, assembly and disassembly. Therefore, akin to the way humans efficiently learn about object and environment physics properties, this work proposes to actively identify system parameters that enable high-precision simulation of real-world manipulation dynamics for volumetric elastoplastic objects.
@@ 2.2 Deformable object modelling
Simulators are essential for advancing robotic manipulation, providing a fast, low-cost alternative to real-world testing (Featherstone, 2014; Collins et al., 2021). In recent years, arguably the most efficient and accurate simulation for 3D deformable objects is achieved by the material point method (MPM) (Jiang et al., 2016). Like most well-known methods for simulating continuum matter, such as position-based dynamics (PBD) and smoothed particle hydrodynamics (SPH) (Yin et al., 2021), MPM represents the object as Lagrangian particles and keeps track of their positions and velocities. Unlike pure Lagrangian methods, MPM achieves faster simulation by computing the motions, deformation gradients, and frictional contacts on a background Eulerian grid (akin to the finite element method), governed by elastic and plastic energy functions and Newton’s laws in the form of partial differential equations (Stomakhin et al., 2013; Jiang et al., 2016; Gao et al., 2017; Hu et al., 2018). MPM has been proven superior to other methods in terms of efficiency and visual effects for objects that undergo large deformations, fractures, and self-collision (Hu et al., 2018). Despite its advantages, accurately identifying physics parameters for real-world robotic manipulation of elastoplastic objects remains an open challenge.
@@ 2.3 Deformable object system identification
Unlike rigid bodies (Jaques et al., 2022; Heiden et al., 2022; Chen et al., 2022), identifying the physics properties of deformable objects is more complex than for rigid bodies, primarily due to the difficulty in measuring key parameters like material properties and friction coefficients (Arriola-Rios et al., 2020; Yin et al., 2021). Previous research has mostly focused on linear and planar deformable objects, such as ropes and cloths (Sundaresan et al., 2022; Yang et al., 2022; Caporali et al., 2024).
Early works that sought to identify 3D volumetric material properties have much simpler assumptions and scenarios. For instance, one of the earlier works uses gradient-based optimisation to retrieve the stiffness of a spring system representing elastic deformable objects (Lloyd et al., 2007), while another work uses an exhaustive search method to find the value of Poisson’s ratio for an elastic form object (Güler et al., 2017). They focus on single parameter identification for volumetric elastic deformable objects with reduced Dofs that under-represent the geometries and deformation behaviours of real-life objects.
Recent efforts like GradSim (Murthy et al., 2020) have used differentiable physics and rendering to identify five parameters for elastic objects with much higher DoFs from single-view simulation videos, demonstrating the feasibility of differentiable system identification with synthetic videos. Two following works, PAC-NeRF (Li et al., 2023) and LPO (Kaneko, 2024) propose to jointly reconstruct object geometries (shapes, positions and colours) and physical properties by using a voxel neural radiance field (Sun et al., 2022) that performs differentiable rendering and allows gradients to be back-propagated from the image space to the Eulerian grid.
Compared to these works that focused on simulations, our study tackles the more challenging task of system identification for real-world objects through robot interactions. Similarly, we also employ the material point method (Hu et al., 2018) and DiffTaiChi (Hu et al., 2020) for differentiable physics simulation and study the system identification task without known object geometries. However, we aim to align simulations with real-world dynamics with minimal robot interactions, using only incomplete, occluded, and noisy point cloud data.
@@ 3.3 Differentiable dynamics modelling
We modified the standard procedures in one simulation step of the MLS-MPM as shown below to incorporate end-effector control following Xian et al. (2023). The readers are referred to the original paper for more details about MPM (Stomakhin et al., 2013) and MLS-MPM (Hu et al., 2018).
@@
    Particle to grid. Use the affine particle-in-cell transform (Jiang et al., 2015) to transfer the velocities and masses of the Lagrangian particles to the background Eulerian grid nodes. In our implementation, we assume equal volume and mass for the particles.
@@
Elastoplasticity: We assume strain elastoplasticity for the studied objects, meaning that the strain-stress relationship of the material is described using the deformation gradient that can be decomposed into elastic and plastic parts , $\mathbf{F}=\mathbf{F}^{E}\mathbf{F}^{P}$, and the first Piola-Kirchoff stress, $\mathbf{P}$, as $\mathbf{P}=\partial\Psi/\partial\mathbf{F}^{E}$, where $\Psi(\mathbf{F}^{E})$ is the elastic energy density function. In this study, we use the fixed corotated model (Stomakhin et al., 2012) as the energy function for its robustness under large deformation, although other constitutive models may be used for different needs. The fixed corotated elastic energy density function is
@@
where $\mathbf{R}^{E}$ is the rotation matrix that comes from the polar decomposition of the deformation gradient, $\mathbf{F}^{E}=\mathbf{R}^{E}\mathbf{S}^{E}$, where $\mathbf{S}^{E}$ is a symmetric matrix. In practice, they are computed from the singular value decomposition of $\mathbf{F}^{E}$ as $\mathbf{F}^{E}=\mathbf{U}^{E}\mathbf{\Sigma}^{E}(\mathbf{V}^{E})^{T}=\mathbf{U}^{E}(\mathbf{V}^{E})^{T}\mathbf{V}^{E}\mathbf{\Sigma}^{E}(\mathbf{V}^{E})^{T}$, and then $\mathbf{R}^{E}=\mathbf{U}^{E}(\mathbf{V}^{E})^{T}$ and $\mathbf{S}^{E}=\mathbf{V}^{E}\mathbf{\Sigma}^{E}(\mathbf{V}^{E})^{T}$ (Baker, 2005).
@@
In this work, we follow Gao et al. (2017) to use the von Mises model (Jones, 2009) to compute the return mapping, which takes on the associative plastic flow assumption. The projection process of the trial stress outside of the yield region can be described succinctly as $\tilde{\boldsymbol{\epsilon}}^{E,n}-\boldsymbol{\epsilon}^{E,n}=\delta_{\gamma}\tilde{\boldsymbol{s}}^{E,n}/||\tilde{\boldsymbol{s}}^{E,n}||$, where $\tilde{\boldsymbol{\epsilon}}^{E,n}=\log(\tilde{\mathbf{F}}^{E,n})$ and $\boldsymbol{\epsilon}^{E,n}=\log(\mathbf{F}^{E,n})$ are the trial and modified Hencky strains, $\tilde{\boldsymbol{s}}^{E,n}=\tilde{\boldsymbol{\epsilon}}^{E,n}-\ tr(\tilde{\boldsymbol{\epsilon}}^{E,n})/3\ \boldsymbol{I}$ is the trial deviatoric stress, and $\delta_{\gamma}\tilde{\boldsymbol{s}}^{E,n}/||\tilde{\boldsymbol{s}}^{E,n}||$ is the modifying term based on the solution given in Gao et al. (2017). In particular, based on the von Mises yield criterion, $\sqrt{3J_{2}}-\boldsymbol{\sigma}_{y}\leq 0$, where $\boldsymbol{\sigma}_{y}$ is the yield stress and $J_{2}$ is the second deviatoric stress invariant, it was derived that $\delta_{\gamma}=||\tilde{\boldsymbol{s}}^{E,n}||-\boldsymbol{\sigma}_{y}/2\mu$. With the fact that the singular vectors of the trial elastic deformation gradient do not change, the return mapping operation is done on the eigenvalues of the trial Hencky stress, thus on the eigenvalues of the trial elastic deformation gradient. Then, we can obtain the result of the return mapping by taking the exponential of the modified deviatoric stress: $\Sigma^{E,n}=e^{\boldsymbol{\epsilon}^{E,n}}$ and then the modified deformation gradient: $\mathbf{F}^{E,n}=\tilde{\mathbf{U}}^{E,n}\mathbf{\Sigma}^{E}(\tilde{\mathbf{V}}^{E,n})^{T}$.
@@
Differentiable programming: Several programming tools are available for creating differentiable simulations, such as PyTorch (Sundaresan et al., 2022; Arnavaz et al., 2023), DiffTaiChi (Hu et al., 2020), Jax (Schoenholz and Cubuk, 2020) and neural-network-based simulators (Heiden et al., 2021). We build our simulator based on DiffTaiChi due to its automated differentiation mechanism, GPU-accelerated parallel computation, fast computation kernel evaluation, intuitive Python APIs, rich community support and various promising applications (Huang et al., 2021; Lin et al., 2022; Xian et al., 2023).
In particular, for each numerical computation step, DiffTaiChi flattens its computation branches (e.g., boundary and collisions) and replaces mutable local variables with extra local storage variables, producing straight-line codes without mutable variables. It then uses standard source code transform to generate the derivative function based on the adjoint method. To compute gradients with a loss function, DiffTaiChi records the order of computation kernels and the scalar variables in the forward simulation direction, and then it computes the gradients for the concerned variables by evaluating the derivative functions in the reversed simulation direction (Hu et al., 2020). We build a simulator based on this programming language to allow automatic gradient computation for the physics parameters. The codes for the simulator will be open-sourced upon paper acceptance.